# GIVT: Generative Infinite-Vocabulary Transformers

### v2

@@ 3.1 VAE training
[…]
Table 1: Results on class-conditional $256{\times}256$ ImageNet, where GIVT-Causal models outperform their quantization-based counterparts. We report FID as well as precision and recall (where available). We use the standard ADM evaluation suite, where FID is calculated w.r.t. the training set. CG: Classifier guidance acceptance rate or scale, CFG $=w$: Classifier-free guidance with weight $w$ [23], DB-CFG $=w$: Our distribution based CFG variant (Sec. 3.3.2), Top-k: Top-k sampling [19] (“mixed” refers to multiple $k$), Beam $=B/F$: Beam search with $B$ beams, $F$ fans, $t$: Temperature sampling by scaling the predicted $\sigma$ of our models with $t$, $t_{C}$: Choice temperature for MaskGIT. Steps shows inference steps. Additional comments: ${}^{\dagger}$umbers obtained by us from public code, ${}^{\star}$nference uses activation caching.
3.2 GIVT training
We next train a GIVT to predict $p(z)$ or $p(z|c)$ (when a conditioning signal $c$ is available, e.g., in class-conditional generation). We flatten $z$ into a $hw$-length sequence of $d$-dimensional real-valued vectors (or “soft tokens”). Note how this differs from the standard VQ-VAE-based setup, where the latent transformer decoder models a $hw$-length sequence of integers denoting codebook indices. Apart from this change, we only perform two small changes to the standard transformer decoder-only architecture (see Fig. 1): We replace the embedding lookup tables at the input with a single linear layer to project from $d$ to the transformer hidden dimension. At the output, we do not predict a categorical distribution, and instead let the transformer predict the parameters of a continuous distribution.
Assuming channel-wise independence, we can model each of the $d$ entries with a $k$-mixture GMM. The GIVT model then predicts $3dk$ parameters ($dk$ means, scales, and mixture probabilities) that we interpret as $d$ scalar GMMs. Experimentally, we found it beneficial to normalize the mixture probabilities with a softmax activation, and the variance parameters with softplus.
We use the standard cross-entropy loss (which is equivalent to the negative log-likelihood) on the distribution $\tilde{p}$ predicted by GIVT, and minimize $\mathcal{L}_{\text{T}}=\sum_{c}\mathbb{E}_{z}\left[-\log\tilde{p}(z|c)\right]$, assuming the the classes or conditioning signal $c$ uniformly distributed (see Appendix A.1 for details on the loss). We train two types of GIVT models, as described next.
GIVT-Causal  Here, $z$ is flattened into a sequence line-by-line, and the GIVT predicts every $d$-dimensional vector in the sequence conditioned on all previous vectors, masking the self-attention layers to be causal [18, 67]. This training strategy is also called teacher forcing and is analogous to the latent modeling in VQ-GAN [18]. For class-conditional image generation we prepend a [CLS] vector to the input sequence, i.e., a learned vector for each class $c$.
GIVT-MaskGIT  Here, as in MaskGIT [6], we mask the input randomly during training and then gradually uncover the masked tokens during inference. The only changes compared to [6] are related to our real-valued tokens: since we have infinitely many tokens, there is no obvious choice to define a special mask token (when using VQ, one can just extend the vocabulary to contain special tokens, such as [MASK]). Instead, given $z$ and a mask $M$ indicating for every spatial location whether it is masked, we first replace the locations in $z$ corresponding to $M$ with zeros (to remove information), and then embed it with a single dense layer, as above. Additionally, we concatenate one of two learned special vectors in the feature dimension, a [MASK] vector for masked locations, and a [UNMASK] vector otherwise (we half the dimension of the embedded inputs and special tokens s.t. the final hidden dimension remains unchanged).
3.3 Inference
Given a VAE and GIVT trained as above, we run GIVT during inference either causally (see Fig. 3) or as in MaskGIT [6]. We now investigate the various inference schemes that have been developed for discrete transformers, and derive their continuous counterparts.
3.3.1 Temperature Sampling, Nucleus Sampling,
Beam Search
In causal sequence models for text (see [24] for an overview and discussion) and VQ-GAN-based approaches, it is common to tune the sampling algorithm. We start with temperature sampling, which for discrete models adapts the softmax temperature of the categorical distributions predicted at each decoding step. For GIVT, we instead scale the standard deviation of the predicted Gaussian distribution(s) and call this strategy “variance scaling”. As we will see in Sec. 4, this simple change can have a significant impact on sample quality.
Nucleus sampling [24] proposes to collect the largest logits such that its cumulative probability after normalization exceeds a threshold (for example 0.8), and to sample from this reduced-support distribution. In GIVT, for a single Gaussian prediction per output dimension, this can be approximated by truncating the predicted distributions per dimension (thereby choosing a higher-density support). We observed that this has a similar effect to variance scaling and therefore do not pursue this strategy.
Figure 4: VAE ablation: We vary the KL weight $\beta$ and number of channels $d$ when training the VAE. Dashed lines show the sampling FIDs obtained with a Base-size GIVT-Causal. As we increase $\beta$, the sampling FID improves, but the reconstruction FID increases, lower bounding the best possible sampling FID. For the main results we use $\beta=0.00005$ and $d=16$.
3.3.2 Distribution-Based Classifier Free Guidance
In the diffusion literature, classifier-free guidance (CFG) [23] has been employed with great success. Concretely, conditional diffusion models are trained with an additional null class $\emptyset$ to learn the unconditional data distribution. Then, during inference, the conditional log density is “moved away” from the unconditional one: given a guidance weight $w$, the updated (diffusion) score estimate is is obtained as
where $\epsilon$ estimates the gradient of the log density of the data distribution, $\epsilon(z,c)\propto\nabla_{z}\log p(z|c)$ (see [23, Sec. 2]). From this, we now derive a CFG variant for our GIVT, since we directly predict a density. We term this approach “Density-Based CFG” (DB-CFG). Eq. 1 can be written as
$\tilde{\epsilon}(z,c)$ $\propto(1+w)\nabla_{z}\log p(z|c)-w\nabla_{z}\log p(z|\emptyset)$
$\propto\nabla_{z}\log\left(p(z|c)^{1+w}p(z|\emptyset)^{-w}\right),$
i.e., $\tilde{\epsilon}$ estimates the log of the density $p_{\text{CFG}}(z|c)\propto p(z|c)^{1+w}p(z|\emptyset)^{-w}$ (see Fig. 5). Thus, we want to adapt our models to sample from $p_{\text{CFG}}$. To this end, we follow [23] and train GIVT with an additional null class $\emptyset$. During inference, we run GIVT twice at every step, once conditional on the actual label $c$ and once conditional on $\emptyset$. To implement classifier-free guidance, we then have to sample from an unnormalized version of $p_{\text{CFG}}(z)$ derived from the two GIVT predictions. To this end, we turn to rejection sampling, which requires: 1) an unnormalized density; 2) a good proposal distribution $p^{\prime}$, that is close to the true target distribution; and 3) a scaling factor $K$ to bound the likelihood between $p^{\prime}$ and the unnormalized target density.
Figure 5: Visualization of our Density-Based Classifier-Free Guidance (DB-CFG). We show the conditional and unconditional PDFs predicted by GIVT, and the resulting CFG PDF for two values of $w$. Note how the CFG distributions become more peaked. We use rejection sampling to sample from $p_{\text{CFG}}$ (see Sec. 3.3.1).
VAE Sample
Step 1
Step 4
Step 8
Step 16
In our setup, the distributions we mix are all Gaussian. We find empirically that the unconditional distributions (i.e., distributions predicted using the $\emptyset$ label) tend to have larger variance than the conditional ones (as visualized in Fig. 5). It is thus sensible to pick sample proposals from $\mathcal{N}(\mu_{c},2\sigma_{c})$, where $\mu_{c},\sigma_{c}$ are the parameters predicted by GIVT when given the label $c$. We empirically find that drawing 1000 samples is enough to find at least one valid sample 99.9% of the time. For the remaining ${<}0.1$%, fall back to sampling from $\mathcal{N}(\mu_{c},\sigma_{c})$.
We emphasize that the overhead of DB-CFG is small: it requires two forward passes (per inference step) instead of one to predict the conditional and unconditional distribution. We then draw 1000 samples from those in parallel on an accelerator, which is very fast. We refer to App. 3.3.2 for Python code.
4 Experiments
We explore various tasks to show the generality of GIVT.
4.1 Image generation
We use ImageNet and explore class-conditional generation (where we condition our GIVT on class labels) for 256px and 512px, as well as unconditional generation for 256px.
VAE  We closely follow the setup of MaskGIT [6]. We use their VAE architecture, built of ResBlocks (as detailed in App. A; encoder and decoder have a combined 53.5M parameters), remove the VQ layer and related losses, and replace it with a linear layer predicting $\mu,\sigma$ (Sec. 3.1). We use the same weights for reconstruction, perceptual, and GAN-loss, as well as identical optimizer parameters, as in [6, 43]; we only vary the latent dimension $d$ and weight $\beta$ of the KL-term. By default, we set the token dimension to 16 (i.e., we predict 16 means and variances per token) and $\beta=5\cdot 10^{-5}$.
GIVT  For GIVT-Causal, we follow the original transformer decoder architecture [67] in decoder-only mode, but remove biases from attention layers, MLP blocks, and LayerNorms, and replace ReLU by GELU as is common practice. For GIVT-MaskGIT, we simply remove the attention mask during training and feed masked inputs instead of shifted ones. Unless otherwise stated, we use the BERT-Large configuration [12]. We configure our GIVT models to predict a single Gaussian distribution by default, but also explore predicting a mixture of 4 distributions per dimension, and predicting a multivariate Gaussian modeling the full covariance matrix of the tokens (instead of assuming a diagonal covariance matrix). For the conditional generation experiments, we use a learned embedding which we prepend to the embedded token sequence. To train GIVT, we use the AdaFactor optimizer from [73] with a cosine schedule (500 epochs; with linear warmup of 50 epochs), set the learning rate and weight decay to $10^{-3}$ and $10^{-4}$, respectively, the optimizer $\beta_{2}$ parameter to $0.95$, the dropout probability to $0.2$ for GIVT-causal and $0.4$ for GIVT-MaskGIT, and the batch size to $8192$. We use the same data augmentation as during VAE training (see [6, 43]), and sample from the VAE encoder distribution for every batch (an additional source of randomness besides data augmentation).
We implement GIVT in JAX [4] and use the JAX version of TensorFlow Probability [14] to implement the predicted distributions and compute the log-probabilities. We train GIVT-Casual and GIVT-MaskGIT models for $256\times 256$ ImageNet, and GIVT-MaskGIT additionally on $512\times 512$. We note that our VAE is trained on $256\times 256$ images, and we use it also for $512\times 512$ experiments (like [6]).
Figure 7: First two plots: Showing the impact of DB-CFG (Sec. 3.3.2) on sampling FID of our class-conditional $256\times 256$ models. First, we show our GIVT-Causal, with three sampling temperatures $t$, and see a compounding effect of $t$ and DB-CFG. Then, we show our GIVT-MaskGIT, using different inference schedules, and observe that apart from a linear schedule, all explored schedules achieve similar quality. Right-most plot: Average standard deviation predicted by GIVT-MaskGIT: when using CFG, we perform inference both with class labels (conditional) and without (unconditional). We plot the average predicted scale over 24 samples for class 94.
GIVT-MaskGIT inference  We fix the number of inference steps to 16 (as in [6]), and explore various inference schedules which define how many tokens are uncovered in each inference step. We employ the cosine schedule from the paper, as well as various power schedules (i.e. letting $r=i/16$ at step $i$, the fraction of masked tokens is given by $\cos(\pi/2r)$ for the cosine schedule, and by $1{-}r^{\alpha}$ for the pow:$\alpha$ schedule). As in [6], we sort tokens by likelihood at each step and sample using a “choice temperature” $t_{C}$.
Exploring the VAE latent space  To better understand the interplay between the feature dimension, the strength of the KL regularization $\beta$, the reconstruction quality of the VAE, and the sampling quality of the GIVT, we train VAEs with latent dimension in $\{4,8,16,32\}$ and $\beta$ in $\{2.5\cdot 10^{-5},5\cdot 10^{-5},10^{-4},2\cdot 10^{-4}\}$ using the VAE-training setup described at the beginning of this section. For each of the resulting VAEs we train a causal GIVT with the smaller BERT-Base [12] dimensions.
Evaluation  For the VAEs we report “reconstruction FID”, the FID obtained when reconstructing the 50k ImageNet validation images. For our GIVT variants and baselines, we report the sampling FID [21] when sampling a balanced set of 50k images covering all ImageNet classes. In both cases, we rely on the well-established ADM TensorFlow Suite [13], which uses the entire ImageNet training set as a reference. Furthermore, we also report Precision and Recall [52]. Finally, we evaluate the representation learning capabilities by training a linear classifier on an average-pooled intermediate representation of unconditional GIVT-Causal as in prior work [8, 71] (see App. D for details).
4.2 Panoptic segmentation and depth estimation
We build on the UViM framework [30], which uses a VQ-VAE to compress the label space of computer-vision dense-prediction tasks, and an encoder-decoder transformer taking the RGB image as an input and predicting the associated dense labels as discrete codes in the VQ-VAE latent space. Here, we replace the VQ-VAE with a $\beta$-VAE and use a GIVT encoder-decoder to model the continuous latent code. For the VAE, we use the same transformer-based autoencoder architecture (6-layer encoder and 12-layer decoder) and cross-entropy loss as [30]. We set $d=16$ and KL weight $\beta=2.5\cdot 10^{-4}$ for panoptic segmentation and $\beta=2\cdot 10^{-4}$ for depth estimation. To build an encoder-decoder GIVT model the same way as in [30], we employ the causal variant described for ImageNet generation and insert a cross-attention layer after each self-attention layer. Following [30], we use the ImageNet-21k-pretrained ViT-L/16 from [56] as the encoder, set the image resolution to 512px, and adopt the preprocessing and optimization hyper-parameters from [30]. We use the UViM variant without encoder and decoder context [30]. Finally, we consider variance scaling and beam search, selecting the best parameters on a held-out subset of the training set as in [30].
5 Results
5.1 Image generation
VAE latent space  In Fig. 4 we show how varying the weight $\beta$ of the KL term affects (i) the VAE reconstruction FID and (ii) the sampling FID of a Base-size GIVT-Causal trained on the corresponding latent sequence. For (i), increasing $\beta$ leads to worse reconstruction FID since the VAE can store less information in the latent. It shifts more of the modeling effort to the VAE decoder, so that the decoder becomes gradually more generative, which affects sampling quality (see [61, Sec., 7] for an extensive discussion).
For (ii), we see the opposite trend: increasing $\beta$ leads to decreased (better) sampling FID for GIVT models trained on the latent sequence. Arguably, this is because the VAE latent sequence more closely follows the Gaussian prior, and hence becomes easier for the GIVT to model. Note, however, that reconstruction FID lower-bounds the sampling FID, so one should chose a VAE with reconstruction FID smaller than the desired sampling FID. For this reason, we chose $\beta=5\cdot 10^{-5}$ for the VAE used in our main models, which obtains a reconstruction FID of 3.66 (note that our best GIVT-Causal model has sampling FID 4.29, see Table 1). Finally, we see that the latent dimension $d$ has quite a small impact on reconstruction and sampling FID.
Sampling FID  In Table 1 we show the sampling FID for three class-conditional model classes on $256\times 256$ ImageNet: GANs, diffusion-based approaches, and sequence modeling approaches. One can observe that our GIVT-Causal models outperform the quantization-based VQ-GAN in terms of FID, even though VQ-GAN has more than 4×4$\times$4 × the model parameters. We study the impact of our DB-CFG for both our models in Fig. 7. On the left, for causal models, we see the impact of temperature scaling, and obtain the best results with $t=0.95$. We did not observe benefits from temperature scaling for our GIVT-MaskGIT models, possibly due to the high-confidence selection-scheme already favoring high-certainty tokens. In the middle of Fig. 7, we instead show the impact of different inference schedules on GIVT-MaskGIT. We find that the cosine schedule used in the original VQ-VAE-based MaskGIT yields the best FID (by a small margin). Moving on to $512\times 512$ ImageNet, we show Sampling FID of available models in Table 2. We see that our GIVT-MaskGIT scales to this resolution, and obtains an FID of 5.81. Finally, we present unconditional results in App. C. This task is considerably harder, but GIVT outperforms the diffusion-based ADM [13] by a solid margin.
@@ 5.1 Image generation
Samples  Fig. 2 shows 10 samples from GIVT-Causal, and App. E shows 10 samples for the GIVT-MaskGIT. We can see that the model produces high-fidelity, coherent samples. To study sample diversity, we sample multiple images for a fixed class in App. E.
@@
In Fig. 6, one can see two samples from our VAE (obtained by decoding latents sampled from the prior), which show a soup of image textures. We then show different steps of the GIVT-MaskGIT inference, and observe similar behavior as in the VQ-based model ([6, Fig. 2]).
[…]
ADM [13] 250 $\numprint{23.2}$0
ADM-G [13] CG $=1.0$ 250 $\numprint{7.72}$
DiT-XL/2 [46] 250 $\numprint{12.03}$
DiT-XL/2-G [46] CFG $=1.5$ 250 $\numprint{3.04}$
[…]
Table 2: Results on class-conditional $512{\times}512$ ImageNet. We use the standard ADM evaluation suite for metrics, where FID is calculated w.r.t. the training set. Our masked GIVT-MaskGIT obtains competitive FID scores with only 16 inference steps. ${}^{\dagger}$umbers obtained by us from public code.
Ablations and visualizations  We did not observe clear benefits from using more than one mixture component per latent dimension in our image generation experiments (both for causal and MaskGIT-based GIVTs). This is perhaps unsurprising since our VAE’s latent codes are regularized towards a unimodal Gaussian prior. Using multiple mixtures might be more beneficial when, e.g., training GIVT on less-regularized latent spaces such as those of vision encoders pretrained in supervised or contrastive settings.
Further, predicting the full covariance matrix per latent vector, rather than assuming a diagonal covariance, only led to very modest gains of about 3% for MaskGIT-based GIVT. On the other hand, a full covariance matrix makes DB-CFG less tractable than the diagonal covariance (because the high dimensional multivariate distribution has more regions of low density).
In Fig. 7 (right), we visualize the predicted standard deviations as a function of the GIVT-MaskGIT inference step. The standard deviation gradually decreases, meaning that the predictions later in the sampling process become more certain. Furthermore, the unconditional predictions generally have a higher standard deviation, as expected.
Representation learning  Table 3 shows the ImageNet linear probing accuracy of unconditional GIVT-Causal and generative models from the literature (we chose the model variants closest in terms of model size and compute). GIVT-Causal matches VIM+ViT (ViT-VQ-GAN) [71] which has more than 2×2$\times$2 × the model parameters and 4×4$\times$4 × the sequence length (and hence FLOPs). GIVT-Causal is only outperformed by MAGE [36], whose latent encoder-decoder architecture is better suited for representation learning than decoder-only models. An investigation of the probing accuracy as function of the layer index can be found in App. D.
@@ 5.2 Panoptic segmentation and depth estimation
Table 4 compares the performance of a GIVT-based UViM variant with a VQ-VAE-based baseline (both without encoder/decoder context) on COCO Panoptic 2017 [29] and NYU Depth v2 [55]. We report the panoptic quality metric (PQ) [29] and RMSE, respectively, and find that our GIVT-based model outperforms the baseline by a small margin in panoptic segmentation and performs slightly worse in depth estimation. In App. E we show visual results.
[…]
recon. inference  recon. inference
UViM [30] 66.0 39.0 0.183 0.459
GIVT (ours) 71.0 40.2 0.195 0.474
@@ 6 Conclusion
In this paper, we proposed two small modifications to standard transformer decoder-only models to make them suitable to generating real-valued vectors. To our knowledge, this is the first decoder-only model amenable to generating sequences of real-valued vectors. In the context of image generation with VQ-GAN or Mask-GIT, this side-steps training difficulties such as low codebook usage in VQ-VAEs and corresponding mitigations like entropy losses or codebook-splitting algorithms, by enabling the use of standard VAEs which are much easier to train. Furthermore, our method avoids large embedding matrices because the feature representations can directly be consumed and predicted by our GIVT model. Surprisingly, our simple, quantization-free approach outperforms its VQ-based equivalent when using causal transformers for class-conditional image generation, and obtains strong performance in dense prediction tasks when applied to UViM. We hope that future work explores applications of GIVT to other modalities such as audio and time-series modeling.

### v4

@@ 5 Results
5.1 Image generation
VAE latent space  In Fig. 5 we show how varying the weight $\beta$ of the KL term affects 1) the VAE reconstruction FID and 2) the sampling FID of a Base-size GIVT-Causal trained on the corresponding latent sequence. For 1), increasing $\beta$ leads to worse reconstruction FID since the VAE can store less information in the latent. It shifts more of the modeling effort to the VAE decoder, so that the decoder becomes gradually more generative, which affects sampling quality (see [67, Sec. 7][55] for more discussion).
For 2), we see the opposite trend: increasing $\beta$ leads to decreased (better) sampling FID for GIVT models trained on the latent sequence. Arguably, this is because the VAE latent sequence more closely follows the Gaussian prior, and hence becomes easier for the GIVT to model. Finally, increasing the number of mixtures $k$ initially reduces the sampling FID substantially, reaching a plateau at $k=16$. We hence set $k=16$ and $\beta=5\cdot 10^{-5}$ by default, and use a VAE with $\beta=10^{-5}$ for the larger (-L) GIVT models. We emphasize that it is common in the literature to explore and tune hyper-parameters such as $\beta$ or analogously the vocab size and commitment loss in VQ [20, Table 5] [55, Table 8][51][46, Fig. 3].
Sampling FID  In Table 1 we show the sampling FID for four model classes on class-conditional $256\times 256$ ImageNet: GANs, diffusion-based approaches, as well as masked and sequence modeling approaches. GIVT-MaskGIT outperforms MaskGIT [7] which has comparable model size and inference cost, and DB-CFG leads to an additional improvement. In absence of guidance techniques, our GIVT-Causal models outperform all diffusion baselines as well as VQGAN by a large margin. Using guidance techniques, GIVT-Causal obtains FID of 3.35 compared to 5.20 for VQGAN with a more than 4.5×4.5$\times$4.5 × smaller model (0.3B for GIVT vs. 1.4B parameters), and also outperforms the 32% larger LDM-4-G. Our larger GIVT-Causal-L+A obtains 16% and 17% reduction in FID without and with guidance, respectively, compared to ViT-VQGAN which has the same generative transformer size but a 4×4$\times$4 × larger sequence length (resulting in more than 4×4$\times$4 × slower sampling) and a 10×10$\times$10 × larger VAE.
We present sampling FID for $512\times 512$ ImageNet in Table 2. GIVT-MaskGIT obtains a 38% lower FID than MaskGIT with comparable model size and inference cost. GIVT-Causal-L+A outperforms DiT-XL/2, the best available DiT model, both without and with guidance (albeit with a larger model).
Finally, we present unconditional results in App. 0.E. This task is considerably harder, but GIVT-Causal beats the diffusion-based ADM [14] by a large margin.
@@ 5.1 Image generation
recon. inference  recon. inference
UViM [33] 66.0 39.0 0.183 0.459
GIVT (ours) 71.0 40.2 0.195 0.474
Ablations and visualizations  Fig. 5 compares the effect of model configuration (number of mixtures $k$, adapter) and sampling algorithm (variance scaling, beam search, DB-CFG) on FID. For every model configuration, all the sampling algorithms lead to solid improvements in FID, with DB-CFG being the most effective one across all configurations. Increasing $k$ from 1 to 16 overall leads to somewhat larger improvements than keeping $k=1$ and adding an adapter. Moreover, combining adapter with $k=16$ results in compounding improvements across sampling algorithms.
Fig. 7 (left) shows the impact of the variance scaling and CFG parameters on the sampling FID. In Fig. 7 (right), we visualize the predicted standard deviation as a function of the GIVT-Causal inference step. The standard deviation gradually decreases, meaning that the predictions later in the sampling process become more certain. Furthermore, the unconditional predictions generally have a higher standard deviation, as expected.
For GIVT-MaskGIT, predicting a single Gaussian with full covariance matrix per latent vector, rather than assuming a diagonal covariance, only led to very modest gains of about 3%. A GMM with factorized component densities therefore seems to be the more effective alternative. Furthermore, a full covariance matrix makes DB-CFG less tractable than the diagonal covariance (because the high dimensional multivariate distribution has more regions of low density).
@@
Samples  Fig. 1 shows ten $512\times 512$ samples from GIVT-Causal-L+A, and App. 0.I shows samples for other GIVT-Causal variants and GIVT-MaskGIT. We can see that the model produces high-fidelity, coherent samples. To study sample diversity, we show multiple samples from different models for a fixed class.
@@
In Fig. 15 in App. 0.I, one can see two samples from our VAE (obtained by decoding latents sampled from the prior), which show a soup of image textures. We then show different steps of the GIVT-MaskGIT inference, and observe similar behavior as in the VQ-based model ([7, Fig. 2]).
@@
Representation learning  Table 4 shows the ImageNet linear probing accuracy of unconditional GIVT-Causal and generative models from the literature (we chose the model variants closest in terms of model size and compute). GIVT-Causal matches VIM+ViT (ViT-VQGAN) [76] which has more than 2×2$\times$2 × the model parameters and 4×4$\times$4 × the sequence length (and hence FLOPs). GIVT-Causal is only outperformed by MAGE [39], whose latent encoder-decoder architecture is better suited for representation learning than decoder-only models. An investigation of the probing accuracy as function of the layer index can be found in App. 0.F.
@@ 5.2 Panoptic segmentation and depth estimation
Table 3 compares the performance of a GIVT-based UViM variant with a VQ-VAE-based baseline (both without encoder/decoder context) on COCO Panoptic 2017 [32] and NYU Depth v2 [61]. We report the panoptic quality metric (PQ) [32] and RMSE, respectively, and find that our GIVT-based model outperforms the baseline in panoptic segmentation and performs slightly worse in depth estimation. In App. 0.I we show visual results.
@@ 6 Conclusion
In this paper, we proposed simple modifications to standard transformer decoder-only models enabling them to generating real-valued vectors. To our knowledge, this is the first decoder-only model amenable to generating sequences of real-valued vectors. In the context of image generation with VQ-GAN or Mask-GIT, this side-steps training difficulties such as low codebook usage in VQ-VAEs and corresponding mitigations like entropy losses or codebook-splitting algorithms, by enabling the use of standard VAEs which are much easier to train. Furthermore, our method avoids large embedding matrices because the feature representations can directly be consumed and predicted by our GIVT model. Our simple, quantization-free approach outperforms its VQ-based counterparts in class-conditional image generation and image representation learning, often by significant margins. GIVT also obtains strong performance in dense prediction tasks when applied to UViM. We hope that future work explores applications of GIVT to other modalities such as audio and time-series modeling.